OpenStudy (anonymous):

Three business partners are to share profits of $24000 in the ratio of 5:4:3. What is the amount of the largest share?

OpenStudy (anonymous):

In order to find how much one portion is, you need to divide the total profit by the number of shares. 5+4+3=? 24,000/?=?

OpenStudy (anonymous):

11 ?

OpenStudy (anonymous):

5+4+3=11 is false,want to try again?

OpenStudy (anonymous):

Oops i ment 12

OpenStudy (anonymous):

Awesome, now that we know this number, we may divide our total profit of 24,000 by 12. 24,000/12=?

OpenStudy (anonymous):

Um 12x2=24 So 200?

OpenStudy (anonymous):

I assume you meant 2,000, but sorta yes. If you take 2,000 and multiply it by the largest share of 5, you will have your answer. 2,000*5=the answer

OpenStudy (anonymous):

Oh so ya gotta do it 2000x5--->2000x4--->2000x3?

OpenStudy (anonymous):

its 10k?

OpenStudy (anonymous):

Yes! Seems like you got it, any questions?

OpenStudy (anonymous):

just fyi can you check like 2 of my questions?

OpenStudy (anonymous):

Sure, just throw 'em at me.

OpenStudy (anonymous):

If an object travels at 22feet per second , how many feet does it travle in 4mins? -- answer 5280ft per second?

OpenStudy (anonymous):

22times 60 x 4

OpenStudy (anonymous):

Yes, you got it! :)

OpenStudy (anonymous):

But take of the per second part.

OpenStudy (anonymous):

:) and the next one is(Peter's weight is 12 pounds more than twice susan weight. find peter weight if together they weigh 150 pounds- is it 104pounds?-- (The next one)-(Half of the people on the bust get off at each stop after the first and no one gets on after the first stop. if only one person got off at stop 5 how many people got on the first stop)

OpenStudy (anonymous):

its 16 the last one.

OpenStudy (anonymous):

104 pounds is correct.

OpenStudy (anonymous):

Sure it's not 32?

OpenStudy (anonymous):

for the last one?

OpenStudy (anonymous):

lets me check

OpenStudy (anonymous):

Yes. 1) 32/2=16 2) 16/2=8 3) 8/2=4 4) 4/2=2 5) 2/2=1

OpenStudy (anonymous):

its 16 or 32?

OpenStudy (anonymous):

Not 100% sure.

OpenStudy (anonymous):

Sorry, I'm sure for all questions except the last.

OpenStudy (anonymous):

i checked it and is it 16? because doing it backwords ya gotta do it x2 so 1x2-2 -- 2x2- 4 4-2-8-16

OpenStudy (anonymous):

Sounds good, I didn't think about going backwords. :) Good job!

OpenStudy (anonymous):

and this one is easy and ------ How many $$$ is for 21 donuts 1donut=0.49$ ------box of 6=1.89$------box of 12=3.59$ = so i just did 3.59+3.59 it will equal 7.18 and i - that by .49 equaling to 6.69

OpenStudy (anonymous):

$6.95 12=3.59 6=1.89 3=1.47 3.59+1.89+1.47=6.95

OpenStudy (anonymous):

12+6+3=21

OpenStudy (anonymous):

How you got that i did 7.18$ and then i subtract it by .49 +6.69

OpenStudy (anonymous):

3.59 for 12, so 3.59*2=7.18 12*2=24 24-21=3 .49*3=1.47 7.18-1.47=5.71 It depends on how the person does the problem.

OpenStudy (anonymous):

.49*21=10.29

OpenStudy (anonymous):

I'm still feeling 6.95 because it is separated accordingly and then added, but it really depends.

OpenStudy (anonymous):

i Think thats incorrect just by saying *(0.49$)*-7.18$)=6.69

OpenStudy (anonymous):

Then go with what you feel is right, seeing as there are multiple "correct" answers.

OpenStudy (anonymous):

Thanks :D for helping me